### v2

# Vision-EKIPL: External Knowledge-Infused Policy Learning for Visual Reasoning

Abstract
Visual reasoning is crucial for understanding complex multimodal data and advancing Artificial General Intelligence. Existing methods enhance the reasoning capability of Multimodal Large Language Models (MLLMs) through Reinforcement Learning (RL) fine-tuning (e.g., GRPO). However, current RL approaches sample action groups solely from the policy model itself, which limits the upper boundary of the model’s reasoning capability and leads to inefficient training. To address these limitations, this paper proposes a novel RL framework called Vision-EKIPL. The core of this framework lies in introducing high-quality actions generated by external auxiliary models during the RL training process to guide the optimization of the policy model. The policy learning with knowledge infusion from external models significantly expands the model’s exploration space, effectively improves the reasoning boundary, and substantially accelerates training convergence speed and efficiency. Experimental results demonstrate that our proposed Vision-EKIPL achieved up to a 5% performance improvement on the Reason-RFT-CoT Benchmark compared to the state-of-the-art (SOTA). It reveals that Vision-EKIPL can overcome the limitations of traditional RL methods, significantly enhance the visual reasoning performance of MLLMs, and provide a new effective paradigm for research in this field.
1 Introduction
Visual reasoning, a core cognitive ability involving interpretation, inference, and logical thinking based on visual information, has emerged as a critical and highly challenging research frontier within the field of Artificial Intelligence [38, 43]. This capability serves as a fundamental cornerstone for numerous complex AI applications, ranging from image recognition [46, 28] and scene understanding [10, 61] to autonomous robotic navigation [29, 36] and autonomous driving [21, 19, 18, 20], underscoring its growing strategic importance.
Figure 1: The output examples of Vision-EKIPL on three visual reasoning tasks, which shows superior reasoning results.
To effectively enhance the visual reasoning capability of machines, the research community has explored diverse technical approaches. Current mainstream research paradigms can be broadly categorized into three types: (1) neural-symbolic methods [13, 4, 8, 67, 16], which aim to integrate the exceptional pattern recognition strengths of deep neural networks with the inherent logical rigor and interpretability of symbolic systems. (2) Supervised Fine-Tuning (SFT) of MLLMs [58, 54], which relies on large-scale annotated datasets for end-to-end training to directly optimize model performance on specific visual reasoning tasks. (3) Reinforcement Learning (RL) based methods [52, 24], exemplified by techniques (e.g., Group Relative Policy Optimization (GRPO) [47]). Such methods leverage RL’s reward mechanisms to activate the latent reasoning potential within pretrained models, demonstrating favorable generalization capability, particularly when tackling complex visual-cognitive tasks involving mathematical logic derivation or code understanding, thus garnering increasing attention.
However, despite the notable successes achieved by RL-based methods on a series of visual reasoning tasks, recent studies [65] have revealed a noteworthy phenomenon: the reasoning paths generated by models post-RL training appear, to a large extent, not to surpass the inherent capability scope of the pretrained foundation model. This suggests that the performance enhancements conferred by RL training might predominantly stem from its role as a preference optimizer. Specifically, RL reinforces the model’s sampling strategy via reward signals, biasing it towards selecting known reasoning paths that have historically yielded high rewards, thereby more efficiently generating correct answers. Yet, this mechanism carries an inherent potential bottleneck: it may excessively favor the exploitation of known successful paths, consequently inhibiting the exploration of novel or more complex reasoning paths. A potential consequence is that the reasoning boundary of a RL-fine-tuned model, compared to its foundation model counterpart with vast potential, might not only fail to expand but could potentially constrict. Furthermore, existing RL methods commonly suffer from slow convergence rates and low training efficiency.
To overcome the dual limitations of current RL methods concerning reasoning boundary expansion and training efficiency, this paper introduces a novel reinforcement learning framework named Vision-EKIPL. Its core innovation lies in significantly broadening the sources of information during the policy learning process. At each input state, the framework not only samples actions based on the current policy model but also incorporates actions from multiple external auxiliary models into the candidate set. Subsequently, these candidate actions are ranked based on the reward signals they receive, and the top-$k$ highest-reward actions are selected to form a high-quality action group. The group is then utilized to guide the optimization of the policy model. Through this mechanism, Vision-EKIPL effectively broadens the policy model’s exploration space by integrating potential solutions offered by diverse "experts" (i.e., the auxiliary models), aiding in the discovery of effective reasoning paths that might be overlooked by a single policy model.
In essence, our method aims to significantly elevate the reasoning frontier of the policy model by proactively introducing and integrating external knowledge (manifested as high-quality actions from auxiliary models) during the optimization process, enabling it to explore and learn richer, more complex reasoning strategies, thereby effectively mitigating the potential reasoning capacity attrition associated with standard reinforcement learning fine-tuning. Concurrently, by directly leveraging high-quality actions from external models to guide the optimization of the policy model, our method also substantially enhances the convergence rate and overall efficiency of the training process. From a broader perspective, this approach can be conceptualized as a promising hybrid paradigm of supervised fine-tuning (data distillation) and reinforcement learning. When the initial reasoning capacity of the policy model is comparatively limited, the model exhibits a greater propensity to select knowledge acquired from external models for supervised learning; conversely, as the policy model’s own capability progressively advances, it gradually leans towards autonomous exploration of deeper reasoning strategies. Vision-EKIPL achieves up to a 5% performance improvement compared to the state-of-the-art(SOTA) on the Reason-RFT-CoT Benchmark. The output examples of Vision-EKIPL are provided in Fig. 1. Although this research primarily validates the efficacy of the framework on visual reasoning tasks, the proposed framework possesses commendable generality and can theoretically be flexibly applied to a broader spectrum of artificial intelligence domains, including various linguistic tasks, visual tasks, and multimodal tasks.
Our main contributions can be summarized as follows:
  • We propose Vision-EKIPL, an innovative reinforcement learning framework that significantly enhances the visual reasoning capability of MLLMs by integrating high-quality actions generated by external models to assist the optimization of the policy model.
  • We demonstrate that incorporating high-quality actions from external models during policy optimization effectively broadens the policy model’s action exploration space, thereby expanding its reasoning boundary.
  • Through extensive experiment evaluation, we verify the effectiveness of the Vision-EKIPL framework, offering valuable insights for advancing visual reasoning research and introducing a new paradigm potentially conducive to promoting multimodal learning research.
2 Related Work
Visual Reasoning
This technology has broad application prospects, including visual counting [38, 37], geometric problem solving [12, 31, 41, 68, 48], visual transformation reasoning [22], scientific research [40, 32], and robotic task planning [23, 29, 17]. Early work in visual reasoning relied on programmatic generation [30, 16, 51] or neuro-symbolic methods [13, 4, 8, 67]. In recent years, driven by the rapid development of MLLMs, the field has seen breakthrough progress. For example, LLaVA-CoT [58] employs a multi-stage Chain-of-Thought (CoT) [57] supervised fine-tuning (SFT) strategy, while Insight-V [11] combines SFT with reinforcement learning (RL). DeepSeek-R1-Zero [14] introduced a rule-based RL approach, significantly enhancing reasoning capability. Building upon DeepSeek-R1 [14], we propose a novel RL method that substantially improves the model’s reasoning performance.
Reinforcement Learning
Reinforcement learning (RL) has demonstrated significant efficacy in enhancing the reasoning capabilities of Large Language Models (LLMs) through iterative, feedback-driven refinement [9, 49, 47, 60, 62, 26, 69]. Notable methodologies include Reinforcement Learning from Human Feedback (RLHF) [44] and Reinforcement Learning from AI Feedback (RLAIF) [6], both of which leverage either human or AI-generated feedback to refine model behavior. Within the domain of vision-language tasks, RL has been successfully employed to align model predictions with human preferences and mitigate the occurrence of hallucinations [50, 63, 64, 70]. More recently, advancements such as DeepSeek-R1-Zero [15] have introduced GRPO [47], a technique that utilizes rule-based rewards to strengthen reasoning abilities without requiring supervised fine-tuning. GRPO has been further adapted for specialized applications, with Visual-RFT [39] employing it for visual grounding and Med-R1 [45] applying it to medical reasoning tasks.Vision-R1 [24] and Reason-RFT [52] adopt a two-stage training paradigm—CoT supervised fine-tuning followed by GRPO-based reinforcement fine-tuning—to enhance the reasoning performance of MLLMs. Distinctly, our Vision-EKIPL is the first to leverage high-quality actions generated by external models to guide policy-model optimization, thereby infusing novel reasoning knowledge and significantly advancing reasoning ability.
3 Method
3.1 Preliminaries
Problem Definition.
Visual reasoning[3, 55, 59] can be formally defined as the task of inferring conclusions or answers by jointly analyzing visual and textual information. Given a visual input $I$ (e.g., images or videos) and an associated textual description or question $T$, the objective is to generate a corresponding answer $A$. This process can be formalized as:
$P:(I,T)\rightarrow A$
where $I\in\mathbb{R}^{H\times W\times C}$ denotes the visual input, characterized by height $H$, width $W$, and the number of channels $C$. The textual input $T$ typically consists of natural language queries or descriptions, while the output $A$ represents the inferred answer, which may be expressed in natural language or structured formats. Through this mapping, visual reasoning models are designed to effectively integrate and interpret multimodal information to perform complex reasoning tasks.
Group Relative Policy Optimization (GRPO).
GRPO[47] presents a novel reinforcement learning framework that has demonstrated strong performance in models such as DeepSeek R1[14]. The fundamental objective of GRPO is to enhance the reasoning capability of model by iteratively refining its policy based on the relative performance of sampled actions within a group.
Figure 2: Overview of the proposed Vision-EKIPL framework. Vision-EKIPL samples high-quality action groups from the action sets of the external models and the policy model based on reward function (RF) evaluation, and then optimizes the policy model using the high-quality action group through the GRPO algorithm.
The process commences with the current policy $\pi_{\theta}$ for a given state $s$. A group of $N$ actions, $\{o_{1},o_{2},\ldots,o_{N}\}$, is sampled from the policy’s output distribution $\pi_{\theta}(o|s)$. Each sampled action $o_{i}$ in this group is subsequently evaluated using a reward function $R(o_{i})$, which quantifies the desirability or effectiveness of the action.
A key element of GRPO is the computation of an advantage score for each action. The advantage $A_{i}$ for the action $o_{i}$ is defined as:
$A_{i}$ $=\frac{R(o_{i})-\operatorname{mean}(\{R(o_{1}),R(o_{2}),\dots,R(o_{N})\})}{\operatorname{std}(\{R(o_{1}),R(o_{2}),\dots,R(o_{N})\})}$ (1)
Actions yielding a positive advantage are considered superior to the group average, while those with a negative advantage are deemed inferior. After computing the advantage $A_{i}$, GRPO evaluates the ratio of the probabilities of each action under the updated policy $\pi_{\theta_{\text{new}}}$ and the previous policy $\pi_{\theta_{\text{old}}}$, denoted as $\text{ratio}_{i}$.
$\text{ratio}_{i}=\pi_{\theta_{\text{new}}}(o_{i}~|~s)~/~{\pi_{\theta_{\text{old}}}(o_{i}~|~s)}$ (2)
The policy model parameters $\theta$ are then updated to increase the likelihood of selecting actions that demonstrated positive advantages and decrease the probability of choosing actions with negative advantages. This update is typically performed using gradient-based optimization method. To mitigate excessive policy updates and enhance training stability, GRPO constrains $\text{ratio}_{i}$ to the interval $1-\delta,1+\delta]$. Moreover, to encourage the learned policy to remain in proximity to the reference distribution $\pi_{\text{ref}}$, a Kullback-Leibler (KL) divergence penalty, weighted by a coefficient $\beta$, is integrated into the optimization objective. Finally, the optimization objective of GRPO can be formulated as follows:
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{s\sim Q,\{o_{i}\}_{i=1}^{N}\sim\pi_{old}}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\min\!\Bigl(\text{ratio}_{i}A_{i},\\ &\quad\text{clip}\left(\text{ratio}_{i},1-\epsilon,1+\epsilon\right)A_{i}\Bigr)-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\Bigg]\end{split}$ (3)
where $Q$ denotes the candidate question set, $\mathbb{D}_{\text{KL}}$ denotes the KL regularization. $\pi_{\text{ref}}$ is typically a frozen pre-trained MLLM. In a nutshell, GRPO aims to maximize the expected advantage, often incorporating this KL divergence as a penalty term.
3.2 External Knowledge-Infused Policy Learning
The overall framework of Vision-EKIPL is illustrated in Fig. 2. Vision-EKIPL is a reinforcement learning framework designed to enhance the visual reasoning capability of MLLMs. The key insight of Vision-EKIPL is leveraging high-quality actions generated by external auxiliary models to guide the optimization of the policy model, thereby infusing novel reasoning knowledge and further expanding the model’s reasoning capacity.
Sampling Action Groups beyond policy model. We introduce a total of $M$ auxiliary models to support the learning process. Although the auxiliary actions are not solely derived from $\pi_{\theta_{\text{old}}}$, as long as the set of actions drawn, for example, from previous policy iterations or a dedicated exploration space, collectively encompasses the entire trajectory space that the current target policy $\pi_{\theta_{\text{new}}}$ can generate, then employing these mixed actions in importance sampling remains theoretically valid, with unbiasedness guaranteed. We have proven it theoretically in the appendix A. Given an input state $s=(x,q)$, where $x$ denotes the visual encoding of the input image and $q$ represents the textual encoding of the question, GRPO first samples a group of actions $\{o^{\theta}_{1},\dots,o^{\theta}_{N}\}$ from the current policy $\pi_{\theta}$. Additionally, for each auxiliary model $\pi^{\phi_{j}}$, it samples a corresponding group of actions $\{o^{\phi_{j}}_{1},\dots,o^{\phi_{j}}_{N}\}$. The sampling process is as follows:
$o^{\theta}_{i}\sim\pi_{\theta}(o\mid x,q),\quad\text{for }i=1,2,\dots,N$ (4)
$\begin{split}o^{\phi_{j}}_{i}\sim\pi^{\phi_{j}}(o\mid x,q),\quad\text{for }i=1,2,\dots,N,\\ \quad\text{for }j=1,2,\dots,M\end{split}$ (5)
All these sampled actions are then combined into a total action group $O$:
$O=\{o^{\theta}_{i}\mid i=1,\dots,N\}\;\cup\;\bigcup_{j=1}^{M}\{o^{\phi_{j}}_{i}\mid i=1,\dots,N\}$ (6)
Reward Calculation. Each sampled action $o_{i}$ is assigned a reward $R(o_{i})$ based on verifiable criteria. In the context of visual reasoning tasks, the reward function $R(o_{i})$[52] integrates two components: a format reward $R_{\mathrm{format}}(o_{i})$ and an accuracy reward $R_{\mathrm{acc}}(o_{i})$. The format reward enforces adherence to a structured response format, while the accuracy reward assesses the correctness of the output, thereby striking a balance between structured reasoning and factual accuracy. The reward function is formally defined as:
$R(o_{i})=R_{\text{format}}(o_{i})+R_{\text{acc}}(o_{i}).$ (7)
The reward calculation follows the criteria outlined below:
  • If the response provides a correct final answer,the model receives an positive accuracy reward. Otherwise, the model receives 0 reward. For the specific definition of accuracy reward, please refer to [52].
  • If the response encloses its reasoning within <think></think> tags and its final answer within <answer></answer> tags, the model receives a format reward of +1. Otherwise, the model receives 0 reward.
Action Selection and Advantage Computation. The actions within the action group $O$ are sorted in descending order based on their reward values, and the top-$G$ actions are selected to form the group of high-quality action $T:\{o_{1},o_{2},\dots,o_{G}\}$, along with their corresponding group of rewards $R:\{r_{1},r_{2},\dots,a_{G}\}$.
The rewards within the sampled reward group $R$ are normalized to compute the relative advantages $\{A_{1},A_{2},\dots,A_{G}\}$, which are computed as shown in Eqn. 1. After computing the relative advantages for the action group $T$, the policy model is updated following Eqn. 3.
4 Experiment
4.1 Experimental Details
In this paper, we employ the Reason-RFT-CoT Dataset [52] to evaluate our method. The experiments are organized into the following three task categories: (1) Visual Counting This task assesses multimodal reasoning by integrating linguistic, visual, and mathematical skills to solve arithmetic problems within 3D block-based scenes. (2) Structure Perception This visual reasoning task requires models to interpret structural information across various mathematical geometries, medical imaging, chart layouts, and architectural designs. (3) Spatial Transformation This spatial-visual reasoning task evaluates a model’s ability to infer single-step or multi-step transformation actions by analyzing initial and final visual states of 3D scenes presented from multiple perspectives (e.g., center, left, right). Each task contains in-domain test set and out-of-domain test set. Specific information can be found in [52].
Implementation Details In our experiments, we utilize Qwen2-VL-2B and Qwen2-VL-7B [56] as policy models. For external models, we selected GPT-4o [27] and Gemini-1.5-Pro [53]. Our implementation is based on the open-source frameworks Open-R1 [25] and vLLM [33] to ensure reproducibility of results and system scalability. For hyperparameters, we employed a cosine learning rate schedule with a peak value of $5\times 10^{-7}$ and adopted the AdamW optimizer to optimize the policy model, setting N and G to 8, with a default KL penalty of $\beta=0.005$. Based on the dataset size, we set the number of epochs for Visual Counting task, Spatial Transformation task, and Structure Perception task to 1, 1, and 5, respectively. To ensure stability and statistical significance of the results, we repeat each major experiment three times under the same setting and report the average accuracy across runs as the final result. All experiments are conducted on a server equipped with 8 A100 GPUs.
API Cost Considerations Incorporating high-quality actions generated by external models such as GPT-4o and Gemini-1.5-Pro introduces non-negligible API costs. These costs include both inference latency (typically 1–3 seconds per query depending on sequence length and model load) and monetary expenses (approximately 0.010.01–0.03 per query, varying with provider pricing tiers). Although these costs are relatively minor in small-scale experiments, they may accumulate significantly in large-scale training or deployment scenarios. In future work, exploring cost-effective open-source alternatives or distillation-based approaches to mitigate reliance on expensive APIs could be a promising direction.
Baselines for comparison To evaluate the performance and generalization capabilities of different training strategies, adhering to the settings in [52], the methods compared in this paper are as follows: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with chain-of-thought (CoT) reasoning. (2) RL-based methods——Reason-RFT-Zero, which applies RL training directly to the base model, Reason-RFT, which first performs supervised learning with partial chain-of-thought (CoT) data before RL training and Vision-EKIPL, which integrates the external auxiliary models into the RL training.
To conduct the comprehensive evaluation, we adopt Qwen2-VL-Instruct [56] as the base model, assessing both its 2B and 7B variants to investigate the influence of model scale. Additionally, the most advanced open-source models [5, 1, 35, 7, 42, 2] and proprietary models [27, 53] are incorporated as baselines to assess the performance of various training paradigms.
Table 1: Results on three visual reasoning tasks. The best results among different training paradigms are highlighted in bold, while the second-best results are underlined. “ID” denotes in-domain test data, and “OOD” denotes out-of-domain test data.
Method Visual Counting Structure Perception Spatial Transformation
Clevr-Math Super-Clevr AVG GeoMath Geometry3k AVG TRANCE TRANCE-L TRANCE-R AVG
ID OOD ID OOD ID OOD
Proprietary Models
GPT-4o-2024-08-06 [27] 68.10 34.31 51.20 50.18 43.49 46.83 42.55 28.67 29.76 35.88
Gemini-1.5-Pro [53] 61.80 37.50 49.65 50.12 48.38 49.45 26.22 18.76 19.88 22.77
Open-Source Models
Qwen2.5-VL-3B-Instruct [5] 75.90 39.30 57.60 36.75 37.44 37.09 8.57 8.26 8.31 8.42
Phi-3.5-Vision-4B-Instruct [1] 21.40 15.20 18.30 36.83 50.25 43.54 7.42 2.45 4.02 5.33
Llava-OneVision-7B [35] 69.70 29.10 49.40 77.63 43.66 60.64 10.00 8.33 8.74 9.27
Qwen2.5-VL-7B-Instruct [5] 74.60 35.20 54.90 44.00 45.61 44.80 19.63 13.12 13.42 16.45
InternVL-2.5-8B [7] 93.50 35.30 64.40 63.00 47.32 51.60 7.19 6.62 6.63 6.91
Llama-3.2-11B-Vision [42] 10.30 9.50 9.90 13.75 20.85 17.30 8.22 8.40 9.03 8.47
Pixtral-12B [2] 42.60 22.90 32.75 30.38 36.09 33.23 7.35 5.03 5.22 6.42
Qwen2VL-2B-Instruct
Zero-Shot 82.40 32.00 57.20 25.86 20.63 23.25 3.78 4.60 4.67 4.35
++ ANS-SFT[52] 96.20 39.20 67.70 51.34 22.50 36.92 77.39 49.24 50.33 58.99
++ CoT-SFT[52] 85.50 46.50 66.00 43.05 25.25 34.15 64.37 43.19 42.86 50.14
++ Reason-RFT-Zero[52] 98.40 44.80 71.60 47.68 32.50 40.09 42.13 34.07 33.41 33.74
++ Reason-RFT[52] 96.80 51.20 74.00 49.03 33.13 41.08 74.61 64.05 64.08 67.58
++ Vision-EKIPL(Ours) 99.10 52.30 75.70 49.70 34.50 42.10 78.23 65.12 65.45 69.60
Qwen2VL-7B-Instruct
Zero-Shot 98.60 42.10 70.35 43.30 43.88 43.59 13.53 12.72 12.78 13.01
++ ANS-SFT[52] 95.00 33.90 64.45 51.34 25.38 38.36 82.19 54.29 54.83 63.77
++ CoT-SFT[52] 87.30 42.40 64.85 50.49 33.00 41.75 81.31 47.90 47.80 59.00
++ Reason-RFT-Zero[52] 99.40 53.00 76.20 55.00 54.75 54.88 67.67 57.20 56.15 56.68
++ Reason-RFT[52] 95.60 51.00 73.30 59.27 49.25 54.26 79.97 59.36 58.61 65.98
++ Vision-EKIPL(Ours) 99.70 53.30 76.50 60.10 56.75 58.42 83.32 62.35 60.47 68.71
4.2 Main Results
Results on In-Domain Tasks
To evaluate the In-Domain (ID) performance of Vision-EKIPL relative to different training paradigms and baseline models across visual reasoning tasks, we conducted extensive training and evaluation on 2B/7B models for three tasks. The results, presented in Tab. 1, indicate the following: (1) Visual Counting RL-based methods consistently outperformed all open-source and proprietary baseline models, as well as SFT-based methods, across both 2B and 7B models, with Vision-EKIPL achieving the best performance among the 7B models; (2) Structure Perception RL-based methods surpassed SFT-based methods in the 7B model, while ANS-SFT demonstrated the best performance in the 2B model. CoT-SFT showed limited improvement, potentially because enforced reasoning supervision hindered cognitive enhancement. Furthermore, Vision-EKIPL in the 7B model outperformed all proprietary models and most open-source models, with the exception of InternVL-2.5-8B [7] and Llava-OneVision-7B [34]; (3) Spatial Transformation Vision-EKIPL achieves the highest performance, surpassing all baseline models. Unlike Reason-RFT, Vision-EKIPL does not require supervised fine-tuning to activate its reasoning capacity, yet still outperforms Reason-RFT. This demonstrates that incorporating high-quality actions from external models can effectively raise the model’s reasoning capacity.
Results on Out-of-Domain Generalization
To validate the out-of-domain (OOD) performance of Vision-EKIPL relative to different training paradigms and baseline models across visual reasoning tasks, we conducted comprehensive experiments on 2B/7B models for three tasks. The results, presented in Tab. 1, reveal the following: (1) Visual Counting RL-based methods demonstrate superior generalization capability compared to SFT-based methods in both 2B and 7B models. Specifically, Vision-EKIPL outperforms ANS-SFT by 13% (2B) and 19% (7B), and also surpasses all open-source and proprietary baselines. Notably, compared to traditional RL methods (e.g., Reason-RFT), Vision-EKIPL significantly expands the model’s reasoning boundary, enabling the model to explore and find correct reasoning paths on complex problems that Reason-RFT could not find. (2) Structure Perception RL-based methods consistently outperform SFT-based methods, with Vision-EKIPL achieving the best results in both 2B and 7B models (8% higher than Reason-RFT on 2B model), while Reason-RFT achieves comparable performance in the 2B model. SFT-based methods shows limited impact, especially in the 7B model; (3) Spatial Transformation RL-based methods surpass SFT-based methods in both 2B and 7B models, while significantly outperforming all baseline models. Vision-EKIPL (2B) exhibits exceptional OOD generalization capability, exceeding GPT-4o [27] by 34% and Gemini-1.5-Pro [53] by 47%. Overall, Vision-EKIPL surpasses all open-source and proprietary baselines, as well as other training methods, demonstrating exceptional performance in visual reasoning generalization capability.
Figure 3: Results of different methods on the Spatial Transformation task across training processes. (a) Evaluation results for 2B model on ID task, (b) Evaluation results for 2B model on OOD task.
4.3 Training Efficiency Evaluation
To demonstrate the data efficiency of Vision-EKIPL during training, we trained all methods on the TRANCE dataset and recorded intermediate and validation results, as illustrated in Fig. 3. Vision-EKIPL demonstrates excellent data efficiency in both in-domain (ID) and out-of-domain (OOD) tasks. The main findings include: (1) On ID tasks, Vision-EKIPL surpasses the performance of Reason-RFT using only 25% of the training data (12,800 samples). Furthermore, when Vision-EKIPL underwent SFT training using the CoT dataset before RL training, following the settings in [52], it achieves 93% of Reason-RFT’s performance using only 12% of the training data. (2) On OOD tasks, Vision-EKIPL achieves the performance of Reason-RFT using only 12% of the data, demonstrating strong generalization capability.
4.4 Analysis on the Sources of Action
As illustrated in Fig.4, we tracked the dynamic evolution of the ratio of actions sampled from the external models and the policy model within the group of actions utilized for parameter updates during the training process of the 2B model on the TRANCE dataset.
Figure 4: Ratio of actions sampled from the external models and the policy model
We can observe that, as the number of training iterations increases, the proportion of actions originating from the external models gradually decreases, while the proportion of actions from the policy model itself progressively increases within the group of actions used for updating the policy model’s parameters.
This phenomenon can be attributed to the initial stages of training: the reasoning capability of the policy model is relatively weaker compared to the auxiliary models (or external models). Consequently, actions sampled from the policy model typically receive lower rewards than those provided by the auxiliary models. To effectively guide the optimization direction of the policy model, we prioritize the selection of actions generated by the auxiliary models.
However, with further model optimization and deeper training, the reasoning capability of the policy model improves significantly, and the rewards obtained from its generated actions also increase accordingly. At this stage, to fully leverage the policy model’s own learning outcomes and accelerate its convergence, we increasingly select actions generated by the policy model to drive the model’s optimization.
Figure 5: Pass@k curves of base model,Reason-RFT and Vision-EKIPL on the Spatial Transformation task. Evaluation results for 2B model (a) and 7B model (b) on ID task.
4.5 Push Forward the Boundary of the Model’s Reasoning Ability
As shown in Fig. 5, we conduct a comparative analysis of the Pass@$K$ scores for Vision-EKIPL, Reason-RFT, and the base model under varying $k$ values. The Pass@$K$ metric reflects the likelihood that a model produces at least one correct answer across $K$ independent samples, and thus serves as a proxy for evaluating the upper bound of the model’s reasoning capability when given sufficient exploration opportunities.
At lower values of $k$, both Vision-EKIPL and Reason-RFT outperform the base model, indicating that reinforcement learning helps guide the model toward more accurate reasoning paths in the early stages. However, a noteworthy phenomenon emerges as $k$ increases: the Pass@$K$ score of Reason-RFT begins to fall behind that of the base model. This suggests that although Reason-RFT improves sample efficiency, it may limit the diversity of the model’s reasoning space by overemphasizing high-reward but familiar reasoning patterns, thus constraining its exploratory capacity.
In contrast, Vision-EKIPL demonstrates a consistently superior performance across all $k$ values. Especially at larger $k$, its Pass@$K$ score significantly surpasses both the base model and Reason-RFT. This clearly highlights Vision-EKIPL’s ability to push the reasoning frontier forward. Its core advantage lies in the integration of high-quality actions generated by external expert models during training. These actions serve as distilled supervision targets, enabling the policy model to learn diverse and novel reasoning strategies that would otherwise be inaccessible through self-sampling.
5 Conclusion
In this paper, we propose Vision-EKIPL, a novel reinforcement learning framework designed to enhance the generalization capability of visual reasoning models. This is achieved by skillfully introducing high-quality actions generated by external auxiliary models to guide the optimization of the policy model during the RL training process. This innovative approach significantly expands the model’s exploration space, enabling the model to effectively surpass traditional reasoning boundary, while also substantially accelerating training convergence speed and overall efficiency. Extensive experiments demonstrate the effectiveness of Vision-EKIPL, providing valuable insights for advancing visual reasoning research and introducing a new paradigm in multimodal learning.
Appendix A Importance Sampling Correction and Its Properties
This section presents a rigorous derivation and properties of importance sampling when samples are drawn from a general proposal distribution $q(o\mid s)$, rather than the reference policy $\pi_{\theta_{\mathrm{old}}}$. We introduce the unbiased IS transformation, the self-normalized IS (SNIS) estimator, bias bounds, the effective sample size (ESS), and explicit formulas for mixture proposals.
A.1 Basic Identity and Correction
Consider the expectation under the reference policy:
$J\;=\;\mathbb{E}_{o\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid s)}\big[g(o)\big],$
where $g(o)$ denotes any integrable function (in our case, often the product of a policy ratio and an advantage term, e.g., $g(o)=\tfrac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{\pi_{\theta_{\mathrm{old}}}(o\mid s)}A(o)$). When samples are drawn from a proposal $q(o\mid s)$, the unbiased IS transformation is obtained via the importance weight
$w(o)\;=\;\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{q(o\mid s)},$
yielding
$J$ $=\mathbb{E}_{o\sim\pi_{\theta_{\mathrm{old}}}}\big[g(o)\big]=\mathbb{E}_{o\sim q}\!\left[\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{q(o\mid s)}\,g(o)\right]=\mathbb{E}_{o\sim q}\big[w(o)\,g(o)\big].$ (8)
For $g(o)=\dfrac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{\pi_{\theta_{\mathrm{old}}}(o\mid s)}A(o)$, Eq. (8) reduces to
$J$ $=\mathbb{E}_{o\sim q}\left[\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{q(o\mid s)}\cdot\frac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{\pi_{\theta_{\mathrm{old}}}(o\mid s)}A(o)\right]=\mathbb{E}_{o\sim q}\left[\frac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{q(o\mid s)}\,A(o)\right].$ (9)
Hence, when using samples from a proposal $q$, the correct unbiased estimator requires the ratio $\tfrac{\pi_{\theta_{\mathrm{new}}}(o)}{q(o)}$, rather than $\tfrac{\pi_{\theta_{\mathrm{new}}}(o)}{\pi_{\theta_{\mathrm{old}}}(o)}$ (unless $q=\pi_{\theta_{\mathrm{old}}}$).
A.2 Sample-Based Estimators
Suppose we draw $o_{1},\dots,o_{n}\sim q$, with weights $w_{i}=w(o_{i})$ and function values $g_{i}=g(o_{i})$.
Unnormalized IS:
$\widehat{J}_{\mathrm{IS}}\;=\;\frac{1}{n}\sum_{i=1}^{n}w_{i}\,g_{i},$
which is unbiased under mild regularity conditions.
Self-Normalized IS (commonly used in practice):
$\widehat{J}_{\mathrm{SNIS}}\;=\;\sum_{i=1}^{n}\tilde{w}_{i}\,g_{i},\qquad\tilde{w}_{i}\;=\;\frac{w_{i}}{\sum_{j=1}^{n}w_{j}}.$
This reduces variance but introduces a small bias that vanishes as $n\to\infty$.
A.3 Effective Sample Size (ESS)
The dispersion of normalized weights can be measured by the effective sample size:
$\mathrm{ESS}\approx\frac{\big(\sum_{i=1}^{n}w_{i}\big)^{2}}{\sum_{i=1}^{n}w_{i}^{2}}.$
When $q$ is close to $\pi_{\theta_{\mathrm{old}}}$, the weights are nearly constant and ESS $\approx n$; if weights are highly imbalanced, ESS becomes much smaller, leading to unstable estimates. In practice, monitoring ESS, weight variance, or high quantiles (e.g., 95% percentile of weights) provides insight into IS quality.
A.4 Bias Bound via Total Variation
If one ignores the correction and directly uses $\mathbb{E}_{q}[g]$ as an approximation to $J$, the bias can be bounded as
$\left|\mathbb{E}_{\pi_{\theta_{\mathrm{old}}}}[g]-\mathbb{E}_{q}[g]\right|$ $=\left|\int g(o)\big(\pi_{\theta_{\mathrm{old}}}(o)-q(o)\big)\,\mathrm{d}o\right|\leq\|g\|_{\infty}\int\big|\pi_{\theta_{\mathrm{old}}}(o)-q(o)\big|\,\mathrm{d}o$
$=2\|g\|_{\infty}\cdot\mathrm{TV}\!\big(\pi_{\theta_{\mathrm{old}}},q\big),$ (10)
where $\mathrm{TV}(p,q)=\tfrac{1}{2}\int|p-q|$ is the total variation distance. Thus, the approximation bias is jointly controlled by the function magnitude and the distributional divergence.
A.5 Explicit Weights for Mixture Proposals
In many settings, the proposal $q$ is a mixture of multiple policies. For instance, if samples are pooled from the reference policy $\pi_{\theta_{\mathrm{old}}}$ and $M$ auxiliary policies $\{\pi_{\phi_{j}}\}_{j=1}^{M}$, the unfiltered proposal can be modeled as
$q(o\mid s)\;=\;\frac{1}{M+1}\Big(\pi_{\theta_{\mathrm{old}}}(o\mid s)+\sum_{j=1}^{M}\pi_{\phi_{j}}(o\mid s)\Big),$
with the importance weight
$w(o)\;=\;\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{\tfrac{1}{M+1}\big(\pi_{\theta_{\mathrm{old}}}(o\mid s)+\sum_{j=1}^{M}\pi_{\phi_{j}}(o\mid s)\big)}.$
If additional selection is performed (e.g., top-$G$ filtering), the actual proposal becomes the conditional distribution
$q_{\text{sel}}(o\mid s)\;=\;\frac{q(o\mid s)\,\mathbf{1}\{o\in S\}}{\int q(o\mid s)\,\mathbf{1}\{o\in S\}\,\mathrm{d}o},$
where $S$ denotes the selection set. Corresponding weights are then computed as $\pi_{\theta_{\mathrm{old}}}(o\mid s)/q_{\text{sel}}(o\mid s)$.
A.6 Algorithmic Pseudocode
Algorithm 1 Importance-Weighted Estimation with Proposal $q$
1:Samples $\{o_{i}\}_{i=1}^{n}\sim q(o\mid s)$; function values $g_{i}=g(o_{i})$; reference and candidate policy densities
2:IS estimate $\widehat{J}_{\mathrm{IS}}$, SNIS estimate $\widehat{J}_{\mathrm{SNIS}}$, ESS
3:Compute weights $w_{i}\leftarrow\dfrac{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid s)}{q(o_{i}\mid s)}$
[…]
5:Normalize: $\tilde{w}_{i}\leftarrow\dfrac{w_{i}}{\sum_{j=1}^{n}w_{j}}$
[…]
8:return $\widehat{J}_{\mathrm{IS}},\ \widehat{J}_{\mathrm{SNIS}},\ \mathrm{ESS}$
A.7 Discussion and Practical Remarks
  • If $q=\pi_{\theta_{\mathrm{old}}}$, then $w\equiv 1$, and Eq. (9) reduces to the standard policy ratio form $\tfrac{\pi_{\theta_{\mathrm{new}}}}{\pi_{\theta_{\mathrm{old}}}}A(o)$.
  • When $q$ deviates substantially from $\pi_{\theta_{\mathrm{old}}}$, IS or SNIS correction is essential. Monitoring ESS and weight statistics provides diagnostics for stability.
  • If $q$ is analytically intractable (e.g., after complex filtering), approximate calculations of $q(o)$ and corresponding bias analysis should be reported. Ablation comparing (A) uncorrected ratios, (B) exact IS, and (C) SNIS can empirically validate the approximation.
A.8 Summary
For samples drawn from a general proposal $q$, unbiased estimation of expectations under $\pi_{\theta_{\mathrm{old}}}$ requires incorporating importance weights $w=\tfrac{\pi_{\theta_{\mathrm{old}}}}{q}$. In the common case $g(o)=\tfrac{\pi_{\theta_{\mathrm{new}}}}{\pi_{\theta_{\mathrm{old}}}}A(o)$, this is equivalent to using the corrected ratio $\tfrac{\pi_{\theta_{\mathrm{new}}}}{q}$. Approximate forms that drop $w$ may still be employed in practice but should be accompanied by theoretical justification (e.g., bias bound in Eq. (10)) and empirical validation via ESS and ablation studies.